What frequency will Johnny hear as a jet, whose engines emit sound at a frequency of 1550 Hz, flies toward her at a speed of 200.0 m/s? Use vS = 343 m/s.

Use the correct units in your answer.



1
Expert's answer
2021-04-09T06:38:38-0400
"f=f_0\frac{v+v_0}{v-v_s}\\f=1550\frac{343+0}{343-200}=3718\ Hz"